# 4DRGS: 4D Radiative Gaussian Splatting for Efficient 3D Vessel Reconstruction from Sparse-View Dynamic DSA Images

### v1

@@ 3.3 Vessel Reconstruction
Given any timestamp $t$, we leverage CUDA-based voxelization in [30] to efficiently query an attenuation volume $\mathbf{V}(t)=\{\mu(\bm{x},t)\}_{\bm{x}\in\mathbf{X}}$. Here, $\mathbf{X}$ is the set of voxel grids defined by the target volume’s resolution and spacing. The final 3D vessel volume is obtained by averaging attenuation volumes across all timestamps $\{t_{j}\}_{j=1}^{T}$ in the complete set of DSA data (Fig. 2(c)): $\overline{\mathbf{V}}=\frac{1}{T}\sum_{j=1}^{T}\mathbf{V}(t_{j})$.
@@ 4.1.2 Implementation Details
DNAF comprises two hash encoders $\bm{h}_{\mathrm{3D}}$ and $\bm{h}_{\mathrm{4D}}$ followed by a two-layer MLP with a width of 64. $\bm{h}_{\mathrm{3D}}$ consists of 12 hash levels. The hash table size is $2^{19}$ with every entry storing a trainable 2-dimensional feature vector. The base resolution is set to 8, increasing by a factor of 1.45 per level. $\bm{h}_{\mathrm{4D}}$ uses a similar configuration, except the base resolution is set to 2, with a growth factor of 1.4. The rasterization and voxelization modules are borrowed from R2-Gaussian [30]. We optimize our model with Adam optimizer [12] for 30k iterations. Adaptive control runs from 600 to 15k iterations, adjusting every 200 iterations with a gradient threshold of 0.0001. The threshold for accumulated attenuation pruning is $\epsilon=1e-6$. A fast version is also provided, where the model trains for 10k iterations, and adaptive control stops at 5k iterations. The learning rates for position, rotation, scale, and DNAF start from 0.0001, 0.001, 0.005, and 0.001, respectively, and exponentially decay to 0.1 of the initial ones. There is also a weight decay factor of $55e-$ 5 for parameters in DNAF. $s_{min}$ and $s_{max}$ are set to 0.1 and 10 times of the voxel spacing, respectively. We initialize $M=30$k kernels with threshold $\delta=0.016$. The loss weight is $\lambda_{ssim}=0.2$.
4.1.3 Evaluation Metrics
@@
We evaluated both 3D vessel reconstruction and 2D DSA image synthesis. For 3D reconstruction, we did not directly compare the reconstructed volumes with the reference volumes due to the unknown data calibration issue. Instead, we evaluated vessel surfaces, which are easier to align and compare. Specifically, we used marching cubes [16] to extract meshes from volumes, with attenuation thresholds of 0.025 for the reference volumes and 0.008 for the algorithm reconstructed volumes. The reference and reconstructed meshes were then aligned using iterative closest point (ICP) [2]. After that, we computed the Chamfer distance (CD) and Hausdorff distance (HD) between the aligned meshes, both measured in millimeters. For 2D image synthesis evaluation, we calculated the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [26] between the synthesized and ground truth DSA images in the test set. We also recorded the running time as an efficiency metric.
4.1.4 Competing Methods
@@
We compared 4DRGS with four methods: FDK with hann filtering [8, 6], 3DGS-based CT reconstruction method R2-Gaussian [30], 3DGS-based DSA image synthesis method TOGS [32], and current SOTA NeRF-based method VPAL [15]. FDK was implemented based on TIGRE-toolbox [3], while others were adapted from their source codes. FDK and R2-Gaussian are static methods, so we directly evaluated their output volumes. TOGS, VPAL, and our 4DRGS are dynamic methods, and we computed the average of generated volumes across timestamps as the final output, as described in Sec. 3.3. TOGS assigns five-element opacity vector to each kernel for dynamic DSA image synthesis and does not support direct volume reconstruction. As a workaround, we uniformly rendered 720 views in a full circle [27] using TOGS and then reconstructed the volume with FDK [6] for each timestamp. All experiments were conducted on a single RTX3090 GPU.
@@ 5 Discussion and Conclusion
5.0.1 Discussion
While our method achieves SOTA results, it has some limitations. For instance, we assume no patient movement during scanning, though minor motion may occur in practice. Additionally, we have not considered calibration errors in scanner geometry and imaging noises in DSA scanning process. We aim to address these limitations in future work to develop a robust clinical-level solution based on our method.
5.0.2 Conclusion
In this work, we present 4DRGS, the first Gaussian splatting-based framework for sparse-view DSA reconstruction. 4DRGS represents vessels using 4D radiative Gaussian kernels, which effectively model static vessel structures and time-varying attenuation changes. We train these kernels with image losses and extract the target vessel volume from them. Two innovations are proposed to enhance reconstruction quality, including accumulated attenuation pruning and bounded scaling activation. Extensive experiments demonstrate the superiority of our method in both 3D vessel reconstruction and 2D DSA image synthesis. Remarkably, our method achieves impressive results within minutes, highlighting its potential for clinical applications.

### v2

@@ 3.3 Vessel Reconstruction
Given any timestamp $t$, we leverage CUDA-based voxelization in [30] to efficiently query an attenuation volume $\mathbf{V}(t)=\{\mu(\bm{x},t)\}_{\bm{x}\in\mathbb{X}}$. Here, $\mathbb{X}\subset\mathbb{R}^{3}$ is the set of voxel grids defined by the target volume’s resolution and spacing. The final 3D vessel volume is obtained by averaging attenuation volumes across all timestamps $\{t_{j}\}_{j=1}^{T}$ in the complete set of DSA data (Fig. 2(c)): $\overline{\mathbf{V}}=\frac{1}{T}\sum_{j=1}^{T}\mathbf{V}(t_{j})$.
@@ 4.1 Experimental Settings
Implementation Details DNAF comprises two hash encoders $\bm{h}_{\mathrm{3D}}$ and $\bm{h}_{\mathrm{4D}}$ followed by a two-layer MLP with a width of 64. $\bm{h}_{\mathrm{3D}}$ consists of 12 hash levels. The hash table size is $2^{19}$ with every entry storing a trainable 2-dimensional feature vector. The base resolution is set to 8, increasing by a factor of 1.45 per level. $\bm{h}_{\mathrm{4D}}$ uses a similar configuration, except the base resolution is set to 2, with a growth factor of 1.4. The rasterization and voxelization modules are borrowed from R2-Gaussian [30]. We optimize our model with Adam optimizer [12] for 30k iterations. Adaptive control runs from 600 to 15k iterations, adjusting every 200 iterations with a gradient threshold of 0.0001. The threshold for accumulated attenuation pruning is $\epsilon=1e-6$. A fast version is also provided, where the model trains for 10k iterations, and adaptive control stops at 5k iterations. The learning rates for position, rotation, scale, and DNAF start from 0.0001, 0.001, 0.005, and 0.001, respectively, and exponentially decay to 0.1 of the initial ones. There is also a weight decay factor of $55e-$ 5 for parameters in DNAF. $s_{min}$ and $s_{max}$ are set to 0.1 and 10 times of the voxel spacing, respectively. We initialize $M=30$k kernels with threshold $\delta=0.016$. The loss weight is $\lambda_{ssim}=0.2$.
@@
Evaluation Metrics We evaluated both 3D vessel reconstruction and 2D DSA image synthesis. For 3D reconstruction, we did not directly compare the reconstructed volumes with the reference volumes due to the unknown data calibration issue. Instead, we evaluated vessel surfaces, which are easier to align and compare. Specifically, we used marching cubes [16] to extract meshes from volumes, with attenuation thresholds of 0.025 for the reference volumes and 0.008 for the algorithm reconstructed volumes. The reference and reconstructed meshes were then aligned using iterative closest point (ICP) [2]. After that, we computed the Chamfer distance (CD) and Hausdorff distance (HD) between the aligned meshes, both measured in millimeters. For 2D image synthesis evaluation, we calculated the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) [26] between the synthesized and ground truth DSA images in the test set. We also recorded the running time as an efficiency metric.
@@
Competing Methods We compared 4DRGS with four methods: FDK with hann filtering [8, 6], 3DGS-based CT reconstruction method R2-Gaussian [30], 3DGS-based DSA image synthesis method TOGS [32], and current SOTA NeRF-based method VPAL [15]. FDK was implemented based on TIGRE-toolbox [3], while others were adapted from their source codes. FDK and R2-Gaussian are static methods, so we directly evaluated their output volumes. TOGS, VPAL, and our 4DRGS are dynamic methods, and we computed the average of generated volumes across timestamps as the final output, as described in Sec. 3.3. TOGS does not support direct volume reconstruction. As a workaround, we uniformly rendered 720 views in a full circle [27] using TOGS and then reconstructed the volume with FDK [6] for each timestamp. All experiments were conducted on a single RTX3090 GPU.
@@ 5 Discussion and Conclusion
Discussion While our method achieves SOTA results, it has some limitations. For instance, we assume no patient movement during scanning, though minor motion may occur in practice. Additionally, we have not considered calibration errors in scanner geometry and imaging noises in DSA scanning process. Moreover, our dataset is a small private set of multi-view cerebral DSA scans. However, single- or bi-plane DSA is more common especially for other body parts like coronary arteries. Our method would fail on such data due to limited views or unmodeled cardiac motion. We leave these aspects for future work.
Conclusion In this work, we present 4DRGS, the first Gaussian splatting-based framework for sparse-view DSA reconstruction. 4DRGS represents vessels using 4D radiative Gaussian kernels, which effectively model static vessel structures and time-varying attenuation changes. We train these kernels with image losses and extract the target vessel volume from them. Two innovations are proposed to enhance reconstruction quality, including accumulated attenuation pruning and bounded scaling activation. Extensive experiments demonstrate the superiority of our method in both 3D vessel reconstruction and 2D DSA image synthesis. Remarkably, our method achieves impressive results within minutes, highlighting its potential for clinical applications.
{credits}
5.0.1 Acknowledgements
This work was supported in part by NSFC grants (No. 6230012077), and HPC Platform of ShanghaiTech University. Shanghai Municipal Central Guided Local Science and Technology Development Fund Project no: YDZX20233100001001.